An automated car wash serves customers with the following serial process: pre-treat, wash, rinse, wax, hand dry. Each of these steps is performed by a dedicated machine except for the hand dry step, which is performed manually on each car by one of 4 workers. The steps of the process have the following processing times 

  • Pre-treat: 5 minute per car 

  • Wash: 3 minutes per car 

  • Rinse: 1 minutes per car 

  • Wax: 4 minutes per car 

  • Hand dry: 12 minutes per car 

a. If the car wash has a demand of 15 cars per hour, what is the flow rate of the process? (Do not round intermediate calculations. Round your answer to 2 decimal places)

b. If the car wash has a demand of 15 cars per hour, what is the utilization of the machine that performs the Wax process? (Round to nearest integer)
Answer #1

a. The capacities of each step is

Pre-treat : 60/5 = 12 per hour

Wash = 60/3 = 20 per hour

Rinse = 60/1 = 60 per hour

Wax = 60/4 = 15 per hour

hand Dry = 60 * 4/ 12 = 20 per hour (* 4 as there are 4 workers)

The bottleneck is Pre treat which is 12 per hour. as demand is higher than that the flow rate will be 12 cars per hour.

b. The wax process is limited by the bottleneck which is Pre-treat @12 cars per hour.

So utilization of Wax machine will be (4 * 12 )/ 60 = 80%